OpenStudy (jazzyfa30):

(6y2 -6/ 8y2 +8y)/(3y -3 / 4y2 + 4y) help please

OpenStudy (jazzyfa30):

@satellite73

OpenStudy (johnweldon1993):

Okay I have a little more time to help you out lol...

OpenStudy (jazzyfa30):

OK THANK YOU

OpenStudy (johnweldon1993):

 Like that??

OpenStudy (jazzyfa30):

yes sorry my internet shut down

OpenStudy (johnweldon1993):

Don't worry about it :) and that is exactly as it is written?

OpenStudy (jazzyfa30):

yes

OpenStudy (johnweldon1993):

Okay...so just like the last one...you flip the bottom fraction and turn this into multiplication.. \[\large \frac{ 6y^2 - 6 }{ 8y^2 + 8y } \times \frac{ 4y^2 + 4y }{ 3y - 3 }\] Now multiply across again...

OpenStudy (jazzyfa30):

9y^3-3=32y^4+32y^2

OpenStudy (johnweldon1993):

Actually you know what....is this a multiple choice question??

OpenStudy (jazzyfa30):

no

OpenStudy (johnweldon1993):

Oh okay...well I would actually just leave it as it....you can't really simplify anything after you set it up like this... \[\large \frac{ 6y^2 - 6 }{ 8y^2 + 8y } \times \frac{ 4y^2 + 4y }{ 3y - 3 }\] this equals... \[\large \frac{ (6y^2 - 6)(4y^2 + 4y) }{ (8y^2 + 8y)(3y - 3) }\] Like I said there's nothing really to simplify so I would leave it at this...if you want to multiply the top and the bottom out...go ahead but there's no reason to in my mind...

OpenStudy (jazzyfa30):

???????

OpenStudy (johnweldon1993):

Whats wrong?

OpenStudy (jazzyfa30):

im confused

OpenStudy (johnweldon1993):

With what part?

OpenStudy (jazzyfa30):

like just multply it out the divide or no

OpenStudy (johnweldon1993):

Well like I said....we can *I'll do it in a second for you* but....there is nothing to do after that to simplify...here let's see...

OpenStudy (johnweldon1993):

\[\large \frac{ (6y^2 - 6)(4y^2 + 4y) }{ (8y^2 + 8y)(3y - 3) }\] After multiplying it out we get... \[\large \frac{ 24y^4 + 24y^3 - 24y^2 - 24y }{ 24y^3 - 24y^2 + 24y^2 - 24y }\] I mean I guess if anything we could factor out a 24y from top and bottom... \[\large \frac{( y^3 + y^2 - y - 1) }{ (y^2 - 1) }\] And this (when divided out) become (y + 1) So I guess I was wrong...good thing you told me to multiply it out lol :) sorry about that...

OpenStudy (jazzyfa30):

so my answer is (y+1)

OpenStudy (johnweldon1993):

That is what I get! :)

OpenStudy (jazzyfa30):

ok what about this one\[\frac{ 1 }{ x-1 }+\frac{ 2 }{ x }=0\]